a 0.30 m radius automobile tire accelerates from rest at a constant 4.0 rad/s^2. what is the centripetal acceleration of a point on the 1/3 radius of the tire after 6.0 sec?
Answer #1

angular velocity after 6 s

w = 4* 6 = 24 rad/s

centripetal acceleration

a = r w^2 = (0.3 / 3)* 24^2

a = 0.24 m/s^2

======

Comment before rate in case any doubt, will reply for sure.. goodluck